OpenStudy (anonymous):

Radical expressions are killing me hoping someone can help me solve some problems!! (9+4i)+(4-3i) (4+4i)-(-8-i)

OpenStudy (amistre64):

the top is just straight addition; get rid of the paranthesis the bottom is trickier; you have to multiply the left side by (-1) and then add everything together

OpenStudy (amistre64):

13 + i is the top 4+4i +8 +i = 12+5i make sense?

OpenStudy (amistre64):

or is this a fraction thing that doesnt have the bar?

OpenStudy (anonymous):

Ok, I got the first one but the second is what is giving me trouble!!

OpenStudy (anonymous):

no 2 separate problems

OpenStudy (amistre64):

radical usually means something under a square root looking sign; rational tends to mean a glorified fraction....

OpenStudy (anonymous):

Yeah there are problems that have the square root sign as well this part of the chapter is called complex numbers

OpenStudy (amistre64):

:) I figured that cause of the "i". The "i" acts like any other variable; but the trick is that when we get i^2 it equals -1 is the only difference .

OpenStudy (anonymous):

Yeah I have seen that in my reading. I wish there were some way to show the nth root of a number because I have some problems with the square root sign and I can't type them on here

OpenStudy (anonymous):

So would the answer to the second one be -12-5i

OpenStudy (amistre64):

I use sqrt(number) and cbrt(number) otherwise its: 4root(number) 6root(number serves the same purpose)

OpenStudy (anonymous):

oh ok, thanks

OpenStudy (amistre64):

if all these equations are are two seperate entities: then the top = 13+i and the bottom = 12+5i

OpenStudy (anonymous):

ok, that second one is confusing!!

OpenStudy (amistre64):

think of the second one like this: -(-8-i) means: -1(-8-i)

OpenStudy (amistre64):

heres a tip: you recall that two negatives equals a positive right?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

then just distribute that (-) thru the brackets to get: --8 --i +8 +i whenever you see two negatives, make them a positive :)

OpenStudy (anonymous):

ok, thanks!!

OpenStudy (amistre64):

for example: 16 -(-4) = 16--4 = 16+4 = 20

OpenStudy (anonymous):

Gotcha!! I think the parenthesis were what was throwing me off for some reason!

OpenStudy (amistre64):

:)